# Measuring the Impact of Missingness in Traffic Stop Data

### v1

@@ Abstract
In this article we explore the data available through the Stanford Open Policing Project. The data consist of information on millions of traffic stops across close to 100 different cities and highway patrols. Using a variety of metrics, we identify that the data is not missing completely at random. Furthermore, we develop ways of quantifying and visualizing missingness trends for different variables across the datasets. We follow up by performing a sensitivity analysis to extend work done on the outcome test as well as to extend work done on sharp bounds on the average treatment effect. We demonstrate that bias calculations can fundamentally shift depending on the assumptions made about the observations for which the race variable has not been recorded. We suggest ways that our missingness sensitivity analysis can be extended to myriad different contexts.
@@ 1 Introduction
Every year, more than 20 million traffic stops are conducted in the United States, representing the most common way in which drivers interact with the police (Davis et al.,, 2018). Unfortunately, not all traffic stops are conducted equally – in a 2019 survey by the Pew Research Center, 59% of Black men and 31% of Black women say that they have been unfairly stopped by the police (Anderson,, 2020). The relationship between belonging to a minority race, such as Black, Latine, and/or Asian Pacific Islander Desi American, and experiencing discriminatory policing is both historical and happening in the present day. Since the late 90’s, concern over racial profiling has led to federal and state mandates requiring the collection of traffic stop data (Russell,, 2001).
The Stanford Open Policing Project (SOPP) [^1] https://openpolicing.stanford.edu/, whose data are the focus of our study, has several dozen datasets representing over 100 million separate traffic stops (Pierson et al.,, 2020). In tandem with the increasing availability of policing data is the growing interdisciplinary scholarship analyzing such data for evidence of racial profiling (Baumgartner et al.,, 2017; Grogger and Ridgeway,, 2006; Pierson et al.,, 2020; Smith and Petrocelli,, 2001; Knox et al.,, 2020). The traffic stops literature develop a range of methods to circumvent the problem of outcome-dependent sampling, as in the outcome test and the veil of darkness; more recent work frames discriminatory policing as a causal inference problem.
@@
Chanin and Welsh’s review of San Diego Police Department traffic stop data (Chanin and Welsh,, 2021) point to the lack of meaningful consideration of data quality in existing traffic stop literature: of the 100 papers they reviewed, the authors find only 19 papers that address missingness and/or recording rates (Chanin and Welsh,, 2021, p. 6). As one may expect, “To an officer who does not see driver race as affecting their own decision-making… data collection only redounds to their detriment” (Chanin and Welsh,, 2021, p. 9).
Motivated by the qualitative evidence of non-ignorable missingness, we conduct a missing data analysis that is organized as follows. Section 3 further contextualizes the missing data problem in the traffic stop setting and defines the different kinds of missingness. Section 3 introduces two metrics for the conditional missingness and numerous visualizations that further diagnose the problem of non-ignorable missingness. Then, we perform two sensitivity analyses in Section 4. The first examines the impact of missing data on the outcome test, a simple descriptive benchmark of the race effect on search from (Pierson et al.,, 2020). The second sensitivity analysis is applied to the causal mediation framework developed by (Knox et al.,, 2020) that bound the Average Treatment Effect (ATE) of race on the use of force during a pedestrian police stop. Our method, summarized in Figure 1, can be applied to datasets and statistics in different domains beyond policing.
@@ 2 Missingness
The problem of missing data is present across multiple disciplines, like healthcare (Kaambwa et al.,, 2012; Ward et al.,, 2020; Hunt et al.,, 2021; Zhang et al.,, 2023), natural disasters (Jones et al.,, 2022), psychology (Gomer and Yuan,, 2023), agriculture (Robbins et al.,, 2013), epidemiology (Allotey and Harel,, 2019; Balzer et al.,, 2020; Tsiampalis and Panagiotakos,, 2020), and crime (Edwards et al.,, 2017; Blasco,, 2021; Stockton et al.,, 2024). Some of them derived methods of imputation (Robbins et al.,, 2013; Blasco,, 2021; Stockton et al.,, 2024; Bounthavong et al.,, 2015) while others provided innovative approaches to combining information across different missingness mechanisms (Ward et al.,, 2020; Gomer and Yuan,, 2023). A few of the papers highlight sensitivity analyses (Allotey and Harel,, 2019; Stockton et al.,, 2024) to obtain a broader view of the missingness landscape, but their work does not use the missing data in the same way as ours does; in particular, they assume the form of the missingness mechanism (e.g., MAR) and we make no such a priori assumptions.
One particular study by Lo et al., (2019) on clinical trial data is closely related to our study. The missingness mechanism in the pharmaceutical setting is “largely related to the post-study reporting of clinical trial data” indicating that, as in traffic stop data, there are incentives to obscure the information. By imputing missing data instead of discarding data with missing values, Lo et al., (2019) achieved superior performance in predicting successful drug approval.
In a pointed summary, Meng, (2024) affirms that the
@@
We expand on the ideas of Meng, (2024) within the context of the SOPP. Part of the motivation for our deep dive into missing data associated with traffic stops is due to policies and laws that require officers to guess at the age, race, and other demographic characteristics of the driver who has been stopped. Guessing at someone’s age or race (and sometimes choosing not to guess) creates complete cases which are systematically different from the incomplete cases (e.g., see Figure 2).
Figure 2 shows two examples of some of the unusual structures in the data: in North Carolina, the age of stopped drivers has peaks in increments of five, which probably reflects some policy of officers recording perceived age instead of actual age. In Seattle, the race for stopped drivers is almost always unrecorded, making us question the fidelity of other variables. The two examples foreshadow the challenge of accurately comparing racial bias across datasets when the data collection schemes are so different. In Section 4, our sensitivity analyses are limited to only datasets where missing race values correspond to non-missing information in the search and contraband found variables.
Figure 2: Examples of idiosyncratic data collection in Charlotte, North Carolina (left) and Seattle, Washington (right). In Charlotte, the age distribution of stopped motorists has peaks at increments of five. In Seattle, the race of motorists is almost never recorded.
@@
  • New York Assembly Bill 3949 (NY3,, )
@@
  • Tennessee House Bill 2167 (TN2,, )
@@
One notable trend of these policies is that usually the race, gender identity, and age is recorded based on officers’ perception. The Stanford Open Policing Project’s own analysis of its data points out this detail (Pierson et al.,, 2020). We can see evidence of “perceived age” in how the age density plots of North Carolina datasets spike at multiples of five (see Figure 2), but we can also see evidence of officer perceptions in the extensive exclusion of both non-binary gender identities and mixed-race identities and the common inclusion of Hispanic, which is an ethnicity, as a level for race. Another important feature of such state mandates are the different variables required for each state. To assess larger trends of missingness, we are able to compare data quality and missing data across datasets with myriad different variables, but datasets that collect neither search nor contraband found (or other acts of force) are outside the scope of most statistical models testing for discriminatory policing.
@@
In a comprehensive assessment of the current missingness literature, Mitra et al., (2023) point out how “literature on causal inference with missing data is very scarce, and it is particularly challenging when both input variables and potential outcomes are systematically missing.” The primary focus of this project is about non-MCAR missingness – diagnosing it and observing how it can impact downstream analyses including causal inference methods. In order to separate these two forms of missingness clearly, we define the following terms.
@@ 2.1.1 Missing data
Missing data refers to an incomplete observed dataset, i.e., data that was intended to have been collected but was not. Missing data is common in all types of data. In the context of policing, because much of the data are self-reported, it is not surprising that there would be extensive missing data. We refer the reader to Meng, (2000), who gives an excellent tour through details of missingness, nonresponse bias, and modeling of real-life missing data mechanisms. In order to describe the missing data mechanisms, we rely on the characterizations from Little and Rubin, (2020):
@@
    If missing data do not depend on the values of the data, missing or observed, that is, the probability of missing is the same for all levels of each variable, the data are called missing completely at random (MCAR) (Little and Rubin,, 2020). Note that the definition implies that the missing data cannot be related to the research question, here, the race effect.
@@
    If the missing data depend on a particular variable only through the observed components of that variable, the missingness mechanism is then called missing at random (MAR) (Little and Rubin,, 2020). Usually, the missingness associated with MAR data has a probability of being missing which is a function of other variables in the dataset which are completely observed.
@@
    We refer to data as Missing Not at Random (MNAR) if the distribution of the missingness depends on the value of the missing components themselves (Little and Rubin,, 2020). The probability of data being missing is related to what the missing values would have been, had we observed them. More recent work has extended MNAR to include any missing data which is not MCAR or MAR, that is, when the function describing the probability of missing data is unknown (Lo et al.,, 2019).
We note that MCAR is ignorable missingness. Estimates computed on MCAR data will typically be more variable (due to smaller sample sizes) but no more biased than estimates computed on complete data. The missingness due to MAR or MNAR is non-ignorable, as estimates will be unreliable and potentially quite biased.
@@ 2.1.2 Missing potential outcomes
While much work has been done to estimate the race effect in traffic stops (see Section 1), the task of identifying race as a causal mechanisms is much harder than estimating the race effect. While certainly race is an individual characteristic, we follow Knox et al., (2020) who consider race to be an aspect of the police-driver encounter rather than an aspect of the driver’s identify. They say, “The manipulation of race is conceptualized as the counterfactual substitution of an individual with a different racial identity into the encounter, while holding the encounter’s objective context—location, time of day, criminal activity, etc.—fixed. In other words, the ‘treatment’ in this case is the entire ‘bundle of sticks’ encapsulating the race of the civilian—including, for example, skin tone, dialect, and clothing” (Knox et al.,, 2020).
@@
In Figure 3, we extend the Directed Acyclic Graph (DAG) from Knox et al., (2020) to illustrate the causal ordering of variables (and missing potential outcomes) in the SOPP datasets along with the missing data aspects of the variable relationships (indicated by $R$ in the graph); the Knox et al., (2020) DAG is provided by the red lines in Figure 3 (note that they include race in $X_{1}$ because their assumption is that all variables are complete). In their work, Knox et al., (2020) assume that any missing data (across all variables) is MCAR. One of their key assumptions is “Mandatory Reporting” explained as “encounters that escalate to the use of force also trigger a reporting requirement and are, therefore, observed in administrative data” (Knox et al.,, 2020). Indeed, the work of Knox et al., (2020) is about the counterfactual of not having been stopped, a missing potential outcome (and not about missing data). We challenge their assumption (of MCAR) and build on their work (of missing potential outcomes) to incorporate data which is missing not completely at random.
@@
Based on the ideas of Moreno-Betancur et al., (2018) (missing data) and Knox et al., (2020) (missing potential outcomes), we create a DAG model which incorporates both missing data and missing potential outcomes in the context of traffic stops (see Figue 3). The variables and their causal ordering are justified in more detail in Appendix B (in particular, Table 6 describes each connecting arrow), but importantly, in our DAG, race is a variable in the set $X_{2}$ (covariates with missing data) which can causally determine not only the traffic stop (mediator $M$) and the outcome ($Y$) (e.g., being searched) but also the missing data aspect of all three variables (e.g., race ($R_{X_{2}}$), stop ($R_{M}$), search ($R_{Y}$)).
@@
The arrows from $X_{2}$ (missing data covariates) to the missingness indicators $R_{X_{2}}$ suggest that the missingness mechanism is MNAR (as described in Section 2.1). Additionally, the DAG in Figure 3 allows for missing data to be affected by the value of the outcome $Y$ (e.g., whether an individual was searched affects missingness in records of their stop). The DAG in Figure 3 is similar to DAG (J) in Moreno-Betancur et al., (2018, Figure 2) who conclude that it is not possible to recover the relationship between the outcome and the treatment in such an MNAR scenario. Because the relationship is unrecoverable, we use a sensitivity analysis to provide an estimate of the range of possible values of our estimates of interest (see Section 4.2).
@@
Figure 3: Extending the DAG in Knox et al., (2020) to incorporate missing data. Bold arrows depict causal mechanisms that lead to missingness. The presented graph is similar to DAG (J) in Moreno-Betancur et al., (2018, Figure 2) suggesting nonrecoverability and necessitating sensitivity analysis. $X_{1}$ represents completely observed covariates like time and day. $X_{2}$ represents covariates with missing information, like age and race. The Knox et al., (2020) DAG is provided by the red lines.
@@ 3.2 Measuring missingness
After selecting a conditioning variable, we partition the conditioning variable into $B$ ordered bins, which we denote as $b\in\{1,2,\ldots,B\}$. We write the corresponding index set of a bin $b$ as $I_{b}$. There is usually a natural choice of bin – if using date as a conditioning variable, $b$ can be on the units of weeks or days; if using time as a conditioning variable, $b$ can be on the unit of hour. Next, for each $b$, we can compute the percent of missingness for observations in that bin $I_{b}$. The percent of missingness can be calculated either for a single sensitive attribute $X_{\cdot j}$ (e.g., the percent of missingness in race on a particular day) or across all $m-1$ variables excluding the conditioning variable (e.g., the percent of missingness on a particular day, aggregated across all variables).
We define these two measurements of conditional missingness. The first is for a single covariate; the second aggregates over $m-1$ covariates. [^3] The $m-1$ in the denominator is because the missingness of the conditioning variable itself is not considered.
@@
The purpose of dCMR is to see the relationship between missing values and some conditioning variable. While we do not formally test the independence of the conditioning variable and the missingness (as one might do if they were formally testing for non-MCAR), we will later compute the maximal correlation between the conditioning variable and the dCMRs.
The maximal correlation measures non-parametric dependence between two variables (Breiman and Friedman,, 1985). Maximal correlation does not assume a linear or monotonic relationship between two variables (unlike the Pearson correlation or Spearman’s rank correlation, respectively). The maximal correlation finds transformations of each variable that maximizes the population-level correlation. While it can be used to formally test for independence, we use maximal correlation as a summary statistic to measure the dependence of two variables. We use the R package acepack which implements the Alternating Conditional Expectations algorithm to compute the maximal correlation (Spector et al.,, 2016).

### v2

@@ Abstract
In this article we explore the data available through the Stanford Open Policing Project. The data consist of information on millions of traffic stops across close to 100 different cities and highway patrols. Using a variety of metrics, we identify that the data is not missing completely at random. Furthermore, we develop ways of quantifying and visualizing missingness trends for different variables across the datasets. Because of the way we have identified the missingness dependence on key variables, imputation is not possible and a sensitivity analysis is required. We performing a sensitivity analysis to extend work done on the outcome test as well as to extend work done on sharp bounds on the average treatment effect. We demonstrate that bias calculations can fundamentally shift depending on the assumptions made about the observations for which the race variable has not been recorded. We suggest ways that our missingness sensitivity analysis can be extended to myriad different contexts.
@@ 1 Introduction
Every year, more than 20 million traffic stops are conducted in the United States, representing the most common way in which drivers interact with the police (davis2018contacts). Unfortunately, not all traffic stops are conducted equally – in a 2019 survey by the Pew Research Center, 59% of Black men and 31% of Black women say that they have been unfairly stopped by the police (anderson_2020). The relationship between belonging to a minority race, such as Black, Latine, and/or Asian Pacific Islander Desi American, and experiencing discriminatory policing is both historical and happening in the present day. Since the late 90’s, concern over racial profiling has led to federal and state mandates requiring the collection of traffic stop data (russell2001racial).
The Stanford Open Policing Project (SOPP) [^1] https://openpolicing.stanford.edu/, whose data are the focus of our study, has several dozen datasets representing over 100 million separate traffic stops (pierson2020large). [^2] After our manuscript was submitted, SOPP substantially expanded their database in July 2025 and included a Python library for accessing the data https://openpolicedata.readthedocs.io/. In tandem with the increasing availability of policing data is the growing interdisciplinary scholarship analyzing such data for evidence of racial profiling (baumgartner2017racial; grogger2006testing; pierson2020large; smith2001racial; knox2020administrative). The traffic stops literature has developed a range of methods designed to measure racial profiling even in the setting of outcome-dependent sampling, as in the outcome test and the veil of darkness; more recent work frames discriminatory policing as a causal inference problem.
@@
Chanin and Welsh’s review of San Diego Police Department traffic stop data (chanin2021) point to the lack of meaningful consideration of data quality in existing traffic stop literature: of the 100 papers they reviewed, the authors find only 19 papers that address missingness and/or recording rates (chanin2021, p. 6). As one may expect, “To an officer who does not see driver race as affecting their own decision-making… data collection only redounds to their detriment” (chanin2021, p. 9).
Motivated by the qualitative evidence of non-ignorable missingness, we conduct a missing data analysis that is organized as follows. Section 2 further contextualizes the missing data problem in the traffic stop setting and defines the different kinds of missingness. Section 3 introduces two metrics for the conditional missingness and numerous visualizations that further diagnose the problem of non-ignorable missingness. Then, we perform two sensitivity analyses in Section 4. The first examines the impact of missing data on the outcome test, a simple descriptive benchmark of the race effect on search from (pierson2020large). The second sensitivity analysis is applied to the causal mediation framework developed by (knox2020administrative) that bounds the Average Treatment Effect (ATE) of race on the use of force during a pedestrian police stop. Our method, summarized in Figure 1, can be applied to datasets and statistics in different domains beyond policing.
@@ 2 Missingness
The problem of missing data is present across multiple disciplines, like healthcare (kaambwa2012; ward2020; hunt2021; zhang2023machine), natural disasters (jones2022human), psychology (gomer2023realistic), agriculture (robbins2013imputation), epidemiology (allotey2019multiple; balzer2020; tsiampalis2020missing), and crime (edwards2017effect; blasco2021missing; stockton2024now). Some of them derived methods of imputation (robbins2013imputation; blasco2021missing; stockton2024now; bounthavong2015) while others provided innovative approaches to combining information across different missingness mechanisms (ward2020; gomer2023realistic). A few of the papers highlight sensitivity analyses (allotey2019multiple; stockton2024now) to obtain a broader view of the missingness landscape, but their work does not use the missing data in the same way as ours does; in particular, they assume the form of the missingness mechanism (e.g., MAR) and we make no such a priori assumptions.
One particular study by lo2019 on clinical trial data is closely related to our study. The missingness mechanism in the pharmaceutical setting is “largely related to the post-study reporting of clinical trial data” indicating that, as in traffic stop data, there are incentives to obscure the information. By imputing missing data instead of discarding data with missing values, lo2019 achieved superior performance in predicting successful drug approval.
In a pointed summary, meng2024 affirms that the
@@
We expand on the ideas of meng2024 within the context of the SOPP. Part of the motivation for our deep dive into missing data associated with traffic stops is due to policies and laws that require officers to guess at the age, race, and other demographic characteristics of the driver who has been stopped. Guessing at someone’s age or race (and sometimes choosing not to guess) creates complete cases which are systematically different from the incomplete cases (e.g., see Figure 2).
Figure 2 shows two examples of some of the unusual structures in the data: in North Carolina, the age of stopped drivers has peaks in increments of five, which probably reflects some policy of officers recording perceived age instead of actual age. In Seattle, the race for stopped drivers is almost always unrecorded. Note that the lack of race data makes the Seattle SOPP dataset inappropriate for an analysis on racial profiling in traffic stops. The issues in North Carolina and Seattle motivate us to think carefully about using the SOPP for large scale analyses on racial profiling in traffic stops. The problems in the dataset may result from data corruption along the way, or alternatively, they may now be resolved in more current downloads of the data. However, like many others, we use the SOPP, and the two examples foreshadow the challenge of accurately comparing racial bias across datasets when the data collection schemes are so different. In Section 4, our sensitivity analyses are limited to only datasets where missing race values correspond to non-missing information in the search and contraband found variables.
Figure 2: Examples of idiosyncratic data collection from the SOPP database in Charlotte, North Carolina (left) and Seattle, Washington (right). In Charlotte, the age distribution of stopped motorists has peaks at increments of five. In Seattle, the race of motorists is almost never recorded.
@@
  • New York Assembly Bill 3949 (NY3949)
@@
  • Tennessee House Bill 2167 (TN2167)
@@
One notable trend of these policies is that usually the race, gender identity, and age is recorded based on officers’ perception. The Stanford Open Policing Project’s own analysis of its data points out this detail (pierson2020large). We can see evidence of “perceived age” in how the age density plots of North Carolina datasets spike at multiples of five (see Figure 2), but we can also see evidence of officer perceptions in the extensive exclusion of both non-binary gender identities and mixed-race identities and the common inclusion of Hispanic, which is an ethnicity, as a level for race. Another important feature of such state mandates are the different variables required for each state. To assess larger trends of missingness, we are able to compare data quality and missing data across datasets with myriad different variables, but datasets that collect neither search nor contraband found (or other acts of force) are outside the scope of most statistical models testing for discriminatory policing.
@@
In a comprehensive assessment of the current missingness literature, mitra2023 point out how “literature on causal inference with missing data is very scarce, and it is particularly challenging when both input variables and potential outcomes are systematically missing.” The primary focus of this project is about non-MCAR missingness – diagnosing it and observing how it can impact downstream analyses including causal inference methods. In order to separate these two forms of missingness clearly, we define the following terms.
@@ 2.1.1 Missing data
Missing data refers to an incomplete observed dataset, i.e., data that was intended to have been collected but was not. Missing data is common in all types of data. In the context of policing, because much of the data are self-reported, it is not surprising that there would be extensive missing data. We refer the reader to meng2000, who gives an excellent tour through details of missingness, nonresponse bias, and modeling of real-life missing data mechanisms. In order to describe the missing data mechanisms, we rely on the characterizations from little2020:
@@
    If missing data do not depend on the values of the data, missing or observed, that is, the probability of missing is the same for all levels of each variable, the data are called missing completely at random (MCAR) (little2020). Note that the definition implies that the missing data cannot be related to the research question, here, the race effect.
@@
    If the missing data depend on a particular variable only through the observed components of that variable, the missingness mechanism is then called missing at random (MAR) (little2020). Usually, the missingness associated with MAR data has a probability of being missing which is a function of other variables in the dataset which are completely observed.
@@
    We refer to data as Missing Not at Random (MNAR) if the distribution of the missingness depends on the value of the missing components themselves (little2020) or an external unrecorded variable. The probability of data being missing is related to what the missing values would have been, had we observed them. More recent work has extended MNAR to include any missing data that is not MCAR or MAR, that is, when the function describing the probability of missing data is unknown (lo2019).
We note that MCAR and MAR are ignorable missingness. Estimates computed on MCAR/MAR data will typically be more variable (due to smaller sample sizes), but no more biased than estimates computed on complete data. Additionally, ignorable missingness allows for the possibility of performing analyses using only the observed data without introducing bias, but they require appropriate methods for analysis. The missingness due to MNAR is non-ignorable, as estimates will be unreliable and potentially quite biased. We emphasize that “without introducing bias” is the key to ignorable missingness (without the phrase, the idea is meaningless because one could always perform an analysis). Importantly, in order for the missing data mechanism to be truly ignorable, the parameter(s) in the model describing the data should have no tie(s) to the parameter(s) that regulate(s) the missing data mechanism (rubin1976).
@@ 2.1.2 Missing potential outcomes
While much work has been done to estimate the race effect in traffic stops (see Section 1), the task of identifying race as a causal mechanisms is much harder than estimating the race effect. While certainly race is an individual characteristic, we follow knox2020administrative who consider race to be an aspect of the police-driver encounter rather than an aspect of the driver’s identity. They say, “The manipulation of race is conceptualized as the counterfactual substitution of an individual with a different racial identity into the encounter, while holding the encounter’s objective context—location, time of day, criminal activity, etc.—fixed. In other words, the ‘treatment’ in this case is the entire ‘bundle of sticks’ encapsulating the race of the civilian—including, for example, skin tone, dialect, and clothing” (knox2020administrative).
@@
In Figure 3, we extend the Directed Acyclic Graph (DAG) from knox2020administrative to illustrate the causal ordering of variables (and missing potential outcomes) in the SOPP datasets along with the missing data aspects of the variable relationships (indicated by $R$ in the graph); the knox2020administrative DAG is provided by the red lines in Figure 3 (note that they include race in $X_{1}$ because their assumption is that all variables are complete). In their work, knox2020administrative assume that any missing data (across all variables) is MCAR. One of their key assumptions is “Mandatory Reporting” explained as “encounters that escalate to the use of force also trigger a reporting requirement and are, therefore, observed in administrative data” (knox2020administrative). Indeed, the work of knox2020administrative is about the counterfactual of not having been stopped, a missing potential outcome (and not about missing data). We challenge their assumption (of MCAR) and build on their work (of missing potential outcomes) to incorporate data which is missing not completely at random.
@@
Based on the ideas of moreno2018canonical (missing data) and knox2020administrative (missing potential outcomes), we create a DAG model which incorporates both missing data and missing potential outcomes in the context of traffic stops (see Figue 3). The variables and their causal ordering are justified in more detail in Appendix B (in particular, Table 6 describes each connecting arrow), but importantly, in our DAG, race is a variable in the set $X_{2}$ (covariates with missing data) which can causally determine not only the traffic stop (mediator $M$) and the outcome ($Y$) (e.g., being searched) but also the missing data aspect of all three variables (e.g., race ($R_{X_{2}}$), stop ($R_{M}$), search ($R_{Y}$)).
@@
The arrows from $X_{2}$ (missing data covariates) to the missingness indicators $R_{X_{2}}$ suggest that the missingness mechanism is MNAR (as described in Section 2.1). Additionally, the DAG in Figure 3 allows for missing data to be affected by the value of the outcome $Y$ (e.g., whether an individual was searched affects missingness in records of their stop). The DAG in Figure 3 is similar to DAG (J) in moreno2018canonical who conclude that it is not possible to recover the relationship between the outcome and the treatment in such an MNAR scenario. They say (our work is analogous to their m-DAG J):
Both m-DAG E and m-DAG J appear plausible. In both cases, the proportion exposed is nonrecoverable and sensitivity analyses would be required… With m-DAG E, the regression-adjusted exposure-outcome association can be unbiasedly estimated using common methods, but sensitivity analyses would be required with m-DAG J.
Because the relationship is unrecoverable, we use a sensitivity analysis to provide an estimate of the range of possible values of our estimates of interest (see Section 4.2).
@@
Figure 3: Extending the DAG in knox2020administrative to incorporate missing data. Bold arrows depict causal mechanisms that lead to missingness. The presented graph is similar to DAG (J) in moreno2018canonical suggesting nonrecoverability and necessitating sensitivity analysis. $X_{1}$ represents completely observed covariates like time and day. $X_{2}$ represents covariates with missing information, like age and race. The knox2020administrative DAG is provided by the red lines.
@@ 3.2 Measuring missingness
After selecting a conditioning variable, we partition the conditioning variable into $B$ ordered bins, which we denote as $b\in\{1,2,\ldots,B\}$. We write the corresponding index set of a bin $b$ as $I_{b}$. There is usually a natural choice of bin – if using date as a conditioning variable, $b$ can be on the units of weeks or days; if using time as a conditioning variable, $b$ can be on the unit of hour. Next, for each $b$, we can compute the percent of missingness for observations in that bin $I_{b}$. The percent of missingness can be calculated either for a single sensitive attribute $X_{\cdot j}$ (e.g., the percent of missingness in race on a particular day) or across all m1m-1 variables excluding the conditioning variable (e.g., the percent of missingness on a particular day, aggregated across all variables).
We define these two measurements of conditional missingness. The first is for a single covariate; the second aggregates over m1m-1 covariates. [^4] The m1m-1 in the denominator is because the missingness of the conditioning variable itself is not considered.
@@
The purpose of dCMR is to see the relationship between missing values and some conditioning variable. While we do not formally test the independence of the conditioning variable and the missingness (as one might do if they were formally testing for non-MCAR), we will later compute the maximal correlation between the conditioning variable and the dCMRs. In the plots below, the calculation of dCMR is done by averaging over the set of variables (when available) given in Table 4 for each dataset.
The maximal correlation measures non-parametric dependence between two variables (breiman1985estimating). Maximal correlation does not assume a linear or monotonic relationship between two variables (unlike the Pearson correlation or Spearman’s rank correlation, respectively). The maximal correlation finds transformations of each variable that maximizes the population-level correlation. While it can be used to formally test for independence, we use maximal correlation as a summary statistic to measure the dependence of two variables. A high value of maximal correlation indicates that there is a (nonlinear) relationship between the dCMR and the variable of interest (see, for example, Figures 4 - 8 where all of the relationships have quite high maximal correlation). We are not able to use standard tests of MAR (e.g., Little’s Test (little1988)) because our data are not normally distributed. The maximal correlation gives us a sense of dependency between missingness and other variables without using a formal test. We use the R package acepack which implements the Alternating Conditional Expectations algorithm to compute the maximal correlation (spector2016package).